# LLM-Forest: Ensemble Learning of LLMs with Graph-Augmented Prompts for Data Imputation

### v3

@@ 4.3 Ablation Study
Impact of the different components in the prompt. We conducted an ablation study on Meta Strategies (M-S) and Correlations and Distributions (C-D)—two additive components designed to enhance performance (Table 2). Meta Strategies aim to inspire the LLM by offering flexible reasoning approaches, leading to improved accuracy for NPHA (65.48% $\rightarrow$ 66.35%) and Gliomas (84.10% $\rightarrow$ 84.41%). Correlations and Distributions provide dataset-specific context, particularly for imbalanced features, and further boost accuracy for NPHA (65.73% $\rightarrow$ 66.35%) and Gliomas (83.55% $\rightarrow$ 84.41%). These results demonstrate that M-S and C-D complement the core prompt components, enhancing the LLM’s reasoning capabilities and contextual understanding.
@@
Table 2: Impact of the different components in the prompt. Results are with % omitted. M-S refers to Meta Strategies, and C-D refers to Correlations and Distributions.
Impact of tree count on LLM-Forest. In this section, we analyze how the number of trees impacts the performance of the LLM-forest. Firstly, as we mentioned in the main results, the ensemble model outperforms a single LLM. Secondly, while adding more trees allows us to provide LLMs with varied perspectives, the imputation results ultimately rely on the information present in the datasets and the LLM’s external knowledge. Given that the pre-trained model for each few-shot LLM remains the same, thus, the information contained within the entire dataset and the external knowledge are both fixed, this creates an upper limit to the LLM’s inference capabilities. As shown in Table 4, beyond three trees, the performance improvements plateau, with results such as only 0.1% improvement for Cancer and 0.31% for Gliomas when using 7 trees compared to using 3 trees. Therefore, using a large number of trees in the forest may not be necessary, as it does not yield substantial gains.
Effectiveness of confidence levels. In LLM prompts, we include instructions for the models to output a confidence level for each imputed value, which we then use as weights when obtaining the ensemble results. We compute the imputation accuracy within different confidence level, the results are shown in Table 3. High-confidence imputations consistently achieve higher accuracy across all datasets which proves that LLMs can evaluate the quality of their own outputs. This self-assessment capability plays a crucial role in improving the final outcomes by leveraging the strengths of each tree while downplaying less certain imputations. We further discuss the impact of confidence-weighted voting in Appendix A.4.
@@ 4.4 Case study
Scalability and time efficiency analysis. To analyze the scalability and efficiency of LLM-Forest, we randomly sample 1,000–5,000 entries from the Diabetes dataset and measure the runtime for neighbor searching and the entire pipeline. As shown in Figure 3, the runtime for neighbor searching with KNN grows sharply, reaching 10,000 seconds for 5,000 records, while our graph-based retrieval method (Graph_ns) remains below 2,000 seconds, demonstrating superior scalability. For the entire pipeline, KNN’s exponential growth remains a bottleneck, whereas LLM-Forest scales efficiently. Its parallelizable architecture ensures computational efficiency for large datasets while maintaining accuracy and robustness in imputation results.
@@ 5 Related Work
With the advancement of LLMs, recent research has expanded their application beyond text-based natural language processing to more structured formats, such as tabular data Lu et al. (2024b); Zhang et al. (2024); Fang et al. (2024). Reasoning on tabular data with LLMs can be approached in two key areas. (1) Global Table Understanding: GPT models face challenges in processing large tables due to several reasons such as the semi-structure of the tabular data Sui et al. (2023), the limited token capacity of LLMs Chen et al. (2024), and abstract or unseen information provided in table cells Ye et al. (2023). These challenges prevent LLMs from fully capturing the table’s structure and content for downstream tasks. To tackle this, several studies such as TaPas Herzig et al. (2020) and Chain-of-TableWang et al. (2024) have proposed advanced table encoding and table-reasoning-chain approaches to help LLMs better understand and reason with tables. (2) Adaptation to Tabular Formats: LLMs, like GPT4, are primarily trained on textual data, limiting their adaptability and effectiveness when applied to the distinct nature of tabular datasets. Recent works Chen et al. (2022); Cheng et al. (2022); Sui et al. (2023); Jiang et al. (2023); Lu et al. (2024a) have proposed prompting and table pre-processing methods incorporating natural language techniques for improved analysis of tabular data. Notably, previous works have focused on downstream tabular reasoning tasks, such as and fact verification. The application of LLMs in tabular data imputation remains significant room for development.
@@
Tabular data imputation is aimed at filling in missing values within structured datasets. Simple statistical methods, such as mean, mode, and median imputation, are widely used but often lead to biased results. Methods like Multiple Imputation by Chained Equations (MICE) Little and Rubin (2019) and MissForest Stekhoven and Bühlmann (2012) improve accuracy by iteratively estimating missing values based on other observed features; However, these methods struggle with high-dimensional and imbalanced data. Recently, deep learning-based models such as GAIN (Generative Adversarial Imputation Networks) Yoon et al. (2018) and DIFFIMPUTE Wen et al. (2024) have been developed to model complex data distributions, which require large and balanced datasets making them less practical for applications where data can be limited. GRAPE You et al. (2020), which encodes relationships between data points by graphs, offers an alternative by leveraging network structures to inform imputation. Leveraging the capabilities of LLMs, Anam Nazir et al. (2023) and CLAIM Hayat and Hasan (2024) explored their use for imputation; however, both approaches require resource-intensive fine-tuning. Our work focuses the less explored area of table imputation with fine-tuning-free approach, marking a new direction within downstream tabular tasks.
@@ 6 Conclusion
In this paper, we propose a novel framework LLM-Forest designed to enhance the data imputation via Large Language Models (LLMs). The framework combines multiple LLM trees employing diverse few-shot learning, forming an LLM forest to achieve robust and accurate imputations. Specifically, to extract high-quality contextual information for LLM trees, we developed a graph-based retrieval algorithm. This algorithm constructs bipartite information graphs, hierarchically merges them to capture cross-feature relationships, and performs random walks on the merged graphs to identify diverse neighboring entries, which are then incorporated into the LLM trees’ prompts. To address the issue of LLM hallucinations, we propose an ensemble approach where multiple LLM trees form an LLM forest, and their outputs are aggregated using confidence-weighted voting. Comprehensive experiments on 9 real-world datasets in various domains demonstrate that LLM-Forest effectively improves imputation accuracy while maintaining scalability and efficiency. The results highlight the potential of combining ensemble learning of LLMs with graph-based retrieval methods to tackle the complex problem of data imputation.
@@ A.2 Experiments Setting
Datasets construction. Our experiments are conducted on 9 real-world public datasets (NPHA, Gliomas, Diabetes, Credit-g, Concrete, Yacht, Wine, Housing and Cancer), and the statistics of the processed datasets are shown in Table 6. These datasets vary in feature types: some consist of categorical features only, others have continuous features only, and a few include a mix of categorical and continuous features. For the Diabetes dataset, we randomly sample 1,000 records for experimentation. To simulate missing data scenarios, we randomly mask 40% of the values in each feature. The datasets are then split into training and testing sets using an 80:20 ratio for downstream classification task evaluation
@@
Baseline methods. We compare our method against widely used tabular data imputation techniques, including approaches such as mode imputation and mean imputation (for numerical datasets only), K-Nearest Neighbors (KNN) Batista et al. (2002), Multiple Imputation by Chained Equations (MICE) Little and Rubin (2019) and Generative Adversarial Imputation Networks (GAIN) Yoon et al. (2018). Since Cancer and Credit-g consist predominantly of non-numerical features, traditional methods like MICE requiring numerical input, cannot be directly applied. For these datasets, we use one-hot encoding to transform non-numerical features into numerical representations before applying imputation. After imputation, we reverse the transformation back to the original feature format and calculate the accuracy of the restored data. Notably, we didn’t include the comparison with CLAIM Hayat and Hasan (2024) since it fine-tunes the LLM for downstream tasks by replacing missing cells with LLM-generated descriptions rather than providing actual imputation values.
Hyperparameter settings. We provide the experiments setting for LLM-Forest and baseline methods. For the GAIN, we used the hyperparameters recommended in the original study. For MICE, we applied the default settings as provided in the PyPI package. For the KNN approach, we selected k = 5 for Gliomas dataset, k = 7 for Daibetes, Cancer and NPHA datasets and k=10 for the rest, as the neighbors we used in LLM-Forest for fair comparison.
For datasets primarily composed of categorical features including NPHA, Gliomas, Diabetes, Cancer and Credit-g, Imputation Accuracy (ACC) is computed based on exact matches between the imputed values and the ground truth. For those numerical features in the categorical datasets, we introduce a small tolerance to account for minor deviations, such as NPHA, Gliomas, and Diabetes, a tolerance of ±$\pm 1$ is used for features like Age and BMI. On Credit-g dataset, for continuous feature credit amount, accuracy is evaluated by matching the most significant digit.
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|x_{i}-\hat{x}_{i}|,$ (9)
@@
Statistic & Deep Learning
@@
Table 7: Comaparison of imputation accuracy of Mixtral-based LLM-Forest with different baseline methods on four datasets with % omitted. The best results are highlighted in boldface. Underlined values indicate the second best.
A.3 Evaluating LLM-Forest with Claude-3.5 and Mixtral
Dataset Mode MICE GAIN KNN LLM-Tree LLM-Forest
MCAR 61.25 57.54 54.18 62.07 62.46 63.18
MAR 63.06 60.75 57.70 63.04 63.03 64.48
MNAR 38.12 41.34 37.37 41.36 49.97 51.25
Dataset Maskn Methods Tree 1 Tree 2 Tree 3 Majority voting Confidence-weighted
@@ A.4 Effectiveness of Graph-based Retrieval and Confidence-weighted Voting
Dataset Maskn Methods Tree 1 Tree 2 Tree 3 Majority voting Confidence-weighted
@@
A.5 Effects on Different Types of Missing Data
In this section, we explore how LLM-Forest  performs under various types of missing data mechanisms.
    The results on Diabetes dataset, presented in Table 8, show that our approach consistently outperforms the baselines across all types of missing data. The largest improvement is observed with MNAR data, where information loss is more significant. In such cases, statistical methods struggle to accurately infer the real relationships between features when entire categories are missing for several features. However, LLMs, leveraging their external knowledge, retain the ability to infer missing categories by understanding the meaning of missing features and analyzing their relationships with existing data, allowing them to perform well even in more challenging scenarios.

### v4

@@ 4.3 Ablation Study
Impact of the different components in the prompt. We conducted an ablation study on Meta Strategies (M-S) and Correlations and Distributions (C-D (Table 2). Meta Strategies aim to inspire the LLM by offering flexible reasoning approaches, leading to improved accuracy for NPHA (65.48% $\rightarrow$ 66.35%) and Gliomas (84.10% $\rightarrow$ 84.41%). Correlations and Distributions provide dataset-specific context, particularly for imbalanced features, and further boost accuracy for NPHA (65.73% $\rightarrow$ 66.35%) and Gliomas (83.55% $\rightarrow$ 84.41%). These demonstrate that M-S and C-D complement the core prompt components, enhancing the LLM’s reasoning capabilities and contextual understanding.
@@
Table 2: Impact of the different components in the prompt. Results are with % omitted. M-S: Meta Strategies, and C-D: Correlations and Distributions.
Impact of tree count on LLM-Forest. We analyze how the number of trees impacts the performance of the LLM-forest. Firstly, as we mentioned in the main results, the ensemble model outperforms a single LLM. Secondly, while adding more trees allows us to provide LLMs with varied perspectives, the imputation results ultimately rely on the information present in the datasets and the LLM’s external knowledge. Given that the pre-trained model for each few-shot LLM remains the same, thus, the information contained within the entire dataset and the external knowledge are both fixed, this poses an upper limit to the LLM’s inference capabilities. As shown in Table 4, beyond three trees, the performance improvements plateau, with results such as only 0.1% improvement for Cancer and 0.31% for Gliomas when using 7 trees compared to using 3 trees. This highlights a trade-off between performance and computational cost. Although more trees can improve accuracy to some extent, the marginal benefits diminish and the additional inference steps introduce overhead. In practice, using a moderate number of trees (e.g., 3 to 5) offers a good balance between efficiency and performance.
Effectiveness of confidence levels. In LLM prompts, we include instructions for the models to output a confidence level for each imputed value, which we then use as weights when obtaining the ensemble results. We compute the imputation accuracy within different confidence level, the results are shown in Table 3. High-confidence imputations consistently achieve higher accuracy across all datasets which proves that LLMs can evaluate the quality of their own outputs. This strong self-assessment capability plays a crucial role in improving the final outcomes by leveraging the strengths of each tree while downplaying less certain imputations. We further discuss the impact of confidence-weighted voting in Appendix A.4, and provide an additional sensitivity analysis of different confidence threshold settings in Appendix A.5.
@@ 4.4 Case study
Scalability and time efficiency analysis. To analyze the scalability and efficiency of LLM-Forest, we randomly sample 1,000–5,000 entries from the Diabetes dataset and measure the runtime for neighbor searching and the entire pipeline. As shown in Figure 4, the runtime for neighbor searching with KNN grows sharply, reaching 10,000 seconds for 5,000 records, while our graph-based retrieval method (Graph_ns) remains below 2,000 seconds, demonstrating superior scalability. Importantly, the constructed graph is shared across all trees in the ensemble, eliminating the need for repeated computation and further reducing overhead. For the entire pipeline, KNN’s exponential growth remains a bottleneck, whereas LLM-Forest scales efficiently. Its parallelizable architecture ensures computational efficiency for large datasets while maintaining accuracy and robustness in imputation results. We also provide results on a large-scale balanced dataset in Appendix A.6, where LLM-Forest demonstrates great performance and scalability.
@@
Effects on Different Types of Missing Data. We explore how LLM-Forest  performs under various types of missing data mechanisms which we introduce in detail in Appendix A.7. The results on Diabetes dataset (Table 5), show that our approach consistently outperforms the baselines across all types of missing data. The largest improvement is observed with MNAR data, where information loss is more significant. In such cases, statistical methods struggle to accurately infer the real relationships between features when entire categories are missing for several features. However, LLMs, leveraging their external knowledge, retain the ability to infer missing categories by understanding the meaning of missing features and analyzing their relationships with existing data, allowing them to perform well even in more challenging scenarios.
Dataset Mode MICE GAIN KNN LLM-Tree LLM-Forest
MCAR 61.25 57.54 54.18 62.07 62.46 63.18
MAR 63.06 60.75 57.70 63.04 63.03 64.48
MNAR 38.12 41.34 37.37 41.36 49.97 51.25
@@ 5 Related Work
Recent researches have extended the use of LLMs beyond natural language tasks to structured formats like tabular data Lu et al. (2024b); Zhang et al. (2024); Fang et al. (2024); Liu et al. (2025). Reasoning on tabular data with LLMs can be approached in two key areas. (1) Global Table Understanding: GPT models face challenges in processing large tables due to several reasons such as the semi-structure of the tabular data Sui et al. (2023), the limited token capacity of LLMs Chen et al. (2024), and abstract or unseen information provided in table cells Ye et al. (2023). These challenges prevent LLMs from fully capturing the table’s structure and content for downstream tasks. To tackle this, several studies such as TaPas Herzig et al. (2020) and Chain-of-Table Wang et al. (2024b) have proposed advanced table encoding and table-reasoning-chain approaches to help LLMs better understand and reason with tables. (2) Adaptation to Tabular Formats: LLMs, like GPT4, are primarily trained on textual data, limiting their adaptability and effectiveness when applied to the distinct nature of tabular datasets. Recent works Chen et al. (2022); Cheng et al. (2022); Sui et al. (2023); Jiang et al. (2023); Lu et al. (2024a); Zou et al. (2025b) have proposed prompting and table pre-processing methods incorporating natural language techniques for improved analysis of tabular data. Notably, previous works have focused on downstream tabular reasoning tasks, such as QA and classification. The application of LLMs in tabular data imputation remains significant room for development.
@@
Tabular data imputation aims to fill in missing values in structured datasets. Simple statistical methods like mean, mode, and median imputation are widely used but often yield biased results. Methods like Multiple Imputation by Chained Equations (MICE) Little and Rubin (2019) and MissForest Stekhoven and Bühlmann (2012) improve accuracy by iteratively estimating missing values based on other observed features, but struggle with high-dimensional or imbalanced data. Recently, deep learning-based models such as GAIN (Generative Adversarial Imputation Networks) Yoon et al. (2018) and DIFFIMPUTE Wen et al. (2024) have been developed to model complex data distributions, which require large and balanced datasets making them less practical for applications where data can be limited. Graphs have also been proved a powerful tool to model structural data Zheng et al. (2024); He et al. (2023); Fu et al. (2024). GRAPE You et al. (2020), which encodes relationships between data points by graphs, offers an alternative by leveraging network structures to inform imputation. Leveraging the capabilities of LLMs, Anam Nazir et al. (2023) and CLAIM Hayat and Hasan (2024) explored their use for imputation; however, both approaches require resource-intensive fine-tuning. Our work focuses the less explored area of table imputation with fine-tuning-free approach, marking a new direction within downstream tabular tasks.
@@ 6 Conclusion
In this paper, we propose a general and scalable framework, LLM-Forest, designed to enhance LLM-based reasoning through ensemble prompting. The framework combines multiple LLM trees employing diverse few-shot learning, forming an LLM forest to achieve robust and accurate imputations. While LLM-Forest is task-agnostic by design, this work focuses on its instantiation for data imputation. Specifically, to extract high-quality contextual information for LLM trees, we developed a graph-based retrieval algorithm. To mitigate LLM bias and uncertainty, we introduce a confidence-weighted voting mechanism that aggregates outputs from all LLM trees based on their self-assessed confidence levels. Comprehensive experiments on 9 real-world datasets in various domains demonstrate that LLM-Forest effectively improves imputation accuracy while maintaining scalability and efficiency. The results highlight the potential of combining ensemble learning of LLMs with graph-based retrieval methods to tackle the complex problem of data imputation.
@@
Acknowledgements
This work is supported by National Science Foundation under Award No. IIS-2117902. The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government.
@@ A.2 Experiments Setting
Datasets construction. Our experiments are conducted on 9 real-world public datasets (NPHA, Gliomas, Diabetes, Credit-g, Concrete, Yacht, Wine, Housing and Cancer), and the statistics of the processed datasets are shown in Table 7. These datasets vary in feature types: some consist of categorical features only, others have continuous features only, and a few include a mix of categorical and continuous features. For the Diabetes dataset, we randomly sample 1,000 records for experimentation. To simulate missing data scenarios, we randomly mask 40% of the values in each feature. The datasets are then split into training and testing sets using an 80:20 ratio for downstream classification task evaluation. To prevent data leakage, imputation is performed only on the training data, and the trained classifier is then applied directly to the (still masked) test set.
@@
Baseline methods. We compare our method against widely used tabular data imputation techniques, including approaches such as mode imputation and mean imputation (for numerical datasets only), K-Nearest Neighbors (KNN) Batista et al. (2002), Multiple Imputation by Chained Equations (MICE) Little and Rubin (2019) and Generative Adversarial Imputation Networks (GAIN) Yoon et al. (2018). We further include advanced methods such as GRAEPE You et al. (2020), which treats the missing data imputation as graph representation learning task; MIRACLE Kyono et al. (2021), which integrates causal regularization into iterative refinement; HyperImpute Jarrett et al. (2022), a hybrid imputer with automatic model selection; TDM Zhao et al. (2023), a transformed distribution matching approach; and Remasker Du et al. (2024), which adopts a masked autoencoding framework. Since Cancer and Credit-g consist predominantly of non-numerical features, traditional methods like MICE requiring numerical input, cannot be directly applied. For these datasets, we use one-hot encoding to transform non-numeric features into numerical representations before applying imputation. After imputation, we reverse the transformation back to the original feature format and calculate the accuracy of the restored data. Notably, we didn’t include the comparison with CLAIM Hayat and Hasan (2024) since it fine-tunes the LLM for downstream tasks by replacing missing cells with LLM-generated descriptions rather than providing actual imputation values.
@@
Statistic &\& Deep Learning
@@
Table 8: Comparison of imputation accuracy of Mixtral-based LLM-Forest with different baseline methods on four datasets with $\%$ omitted. The best results are highlighted in boldface. Underlined values indicate the second best.
Hyperparameter settings. We provide the experimental setting for LLM-Forest and baseline methods. For MICE, we applied the default settings as provided in the PyPI package. For the KNN approach, we selected k = 5 for Gliomas dataset, k = 7 for Daibetes, Cancer and NPHA datasets and k=10 for the rest, as the neighbors we used in LLM-Forest for fair comparison. For all other baselines, we adopt their original implementations and follow the hyperparameters recommended in their respective papers.
For datasets primarily composed of categorical features including NPHA, Gliomas, Diabetes, Cancer and Credit-g, Imputation Accuracy (ACC) is computed based on exact matches between the imputed values and the ground truth. For those numerical features in the categorical datasets, we introduce a small tolerance to account for minor deviations, such as NPHA, Gliomas, and Diabetes, a tolerance of $\pm 1$ is used for features like Age and BMI. On Credit-g dataset, for continuous feature credit amount, accuracy is evaluated by matching the most significant digit.
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|x_{i}-\hat{x}_{i}|,$ (9)
Dataset Mask% n Methods Tree 1 Tree 2 Tree 3 Majority voting Confidence-weighted
@@
Dataset Mask% n Methods Tree 1 Tree 2 Tree 3 Majority voting Confidence-weighted
@@
A.3 Evaluating LLM-Forest with Claude-3.5 and Mixtral
A.5 Confidence Threshold Sensitivity
We conducted a sensitivity analysis on the Cancer dataset using 7 trees, evaluating several alternative confidence threshold settings. Our default scheme assigns weights of 1.0, 0.6, and 0.3 to “High,” “Medium,” and “Low” confidence predictions, respectively, based on the straightforward principle that more confident predictions should exert greater influence.
As shown in Table 11, the overall performance remains stable across reasonable weight configurations, with only minor variations. A key reason for this stability is that high-confidence predictions consistently dominate the results, as shown in Table 12. However, assigning disproportionately high weights to low-confidence predictions (e.g., [0.1, 1.0, 0.6]) results in performance degradation, sometimes even worse than using a single tree. This suggests that amplifying the influence of unreliable outputs can negatively affect the effectiveness of the ensemble.
Confidence threshold
[High, Medium, Low]
[1.0,0.6,0.3] [1.0,0.4,0.2] [1.0,0.3,0.2] [0.1,1.0,0.6]
Cancer 71.93 71.57 71.47 70.80
Table 11: Performance under different confidence threshold combinations (Cancer dataset). Results are reported with % omitted.
Confidence level tree1 tree2 tree3 tree4 tree5 tree6 tree7
High 79.34 78.21 78.99 81.05 78.79 80.25 79.41
Medium 17.99 18.35 18.25 16.85 19.31 16.56 18.24
Low 2.67 3.44 2.77 2.10 1.90 3.19 2.35
Table 12: Confidence level distribution across trees (with % omitted)
Table 12 further confirms that across different trees, our results are highly consistent, with "High" confidence consistently dominating ( 80%). This consistency indicates that regardless of specific values within a reasonable range, the overall decision-making trend remains unchanged and reliable.